### v2

# TongUI: Building Generalized GUI Agents by Learning from Multimodal Web Tutorials

Abstract
Building Graphical User Interface (GUI) agents is a promising research direction, which simulates human interaction with computers or mobile phones to perform diverse GUI tasks. However, a major challenge in developing generalized GUI agents is the lack of sufficient trajectory data across various operating systems and applications, mainly due to the high cost of manual annotations. In this paper, we propose the TongUI framework that builds generalized GUI agents by learning from rich multimodal web tutorials. Concretely, we crawl and process online GUI tutorials (such as videos and articles) into GUI agent trajectory data, through which we produce the GUI-Net-1M dataset containing 1M trajectory data across five operating systems and more than 280 applications. We develop the TongUI agent by fine-tuning Qwen2.5-VL-3B/7B models on GUI-Net-1M, which shows remarkable performance improvements on commonly used grounding and navigation benchmarks, outperforming baseline agents about 10% on multiple benchmarks, showing the effectiveness of the GUI-Net-1M dataset and underscoring the significance of our TongUI framework. We have fully opened-source the code, the GUI-Net-1M dataset, and the trained models.
Figure 1: TongUI crawls and processes multimodal web tutorials from diverse sources into GUI trajectories. The collected GUI-Net-1M dataset spans over 280 applications across five operating systems, leading to powerful generalization in a wide range of GUI tasks.
1 Introduction
Graphical User Interface (GUI) agents based on large foundation models are designed to automate tasks on digital devices by emulating human interactions with a variety of applications on different devices [24, 37, 46]. These agents utilize large language models (LLMs) or vision-language models (VLMs) to process visual inputs (screenshots) and textual inputs (accessibility tree and HTML code) and produce corresponding actions, such as clicking buttons, filling forms, and scrolling, to complete GUI tasks [14, 10, 45]. GUI agents significantly enhance human-computer interaction, providing potential applications to various domains, such as software testing, financial services, office assistance, and industrial automation, improving work efficiency and user experience.
Recently, notable efforts have been made for GUI agents by fine-tuning the foundation models [18, 28, 42]. However, collecting sufficient data for fine-tuning is challenging. Most existing approaches rely on either manually annotated interaction trajectories [20] that are high-quality but costly to obtain, or synthetic data generated from large open-source or proprietary LLM/VLMs, which may lack diversity and accuracy [27, 21, 6]. Desirable GUI agents need to perform actions across a wide variety of applications, each with unique interaction sequences, but collecting comprehensive and diverse data across different applications and operating systems remains a major challenge. As a result, the lack of large-scale, diverse, and well-structured GUI data continues to be a key bottleneck in developing robust and generalized GUI agents.
In this paper, we propose the TongUI framework that builds generalized GUI agents by learning from rich multimodal web tutorials, where ‘Tong’ means generalization in Chinese. Our key observation is that there are readily available multimodal web tutorials on the Internet about how to control computers and smart mobile phones, offering a wealth of information on interacting with various applications and operating systems. These tutorials, in either video or article with screenshots formats, provide detailed step-by-step instructions on interacting with GUI. Compared to the aforementioned manually collected or synthetic data pipeline, online tutorials offer easy accessibility, rich information content, and good quality. Thus, it is natural to build a framework that converts the multimodal web tutorials in diverse formats into task-solving trajectories for GUI agent tuning.
In doing so, we design the TongUI framework composed of four steps: tutorial crawling, tutorial processing, trajectory generation, and data filtering. In the tutorial crawling step, we write some seed tasks and use LLMs to extend them into a wider collection of tasks. The generated tasks serve as keywords for retrieving content from hosts for online GUI tutorials (such as articles from WikiHow and videos from YouTube). The tutorial processing step aims to extract the textual descriptions and screenshots of multimodal tutorials. We first obtain textual descriptions of multimodal tutorials via automatic speech recognition (ASR) or captioning, through which task queries and plans are produced using LLMs on the obtained textual descriptions. Then, we extract salient frames from videos as screenshots of each step, while the images in articles are directly regarded as screenshots. Finally, we leverage one zero-shot GUI agent to automatically recognize trajectories, including reasoning thoughts and actions between two steps. In the data filtering step, we apply a multi-stage pipeline to ensure data quality and GUI relevance, including duplicate tutorial removal, LLM-based content filtering, and trajectory-level filtering using GUI agents. Based on the tutorial collection pipeline, we construct a GUI-Net-1M dataset that contains 1M data across five operating systems with more than 280 applications. Notably, our dataset includes tutorials from different time periods, where evolving layouts still preserve consistent execution logic between similar tasks, helping the model generalize to diverse GUI environments.
Based on GUI-Net-1M, we develop the TongUI agent using Qwen2.5-VL-3B/7B models. We evaluate the TongUI agents on both the offline and online settings, and the results show that the TongUI agent exhibits consistent improvements in grounding and navigation capabilities. The results demonstrate underscores the effectiveness of the TongUI framework and the collected GUI-Net-1M dataset that improve the agents in a low-cost manner without the need for expensive manual annotation.
In summary, our contributions are three-fold:
(1) We propose the TongUI framework that enables GUI agents to automatically learn from web resources via tutorial crawling, tutorial processing, trajectory generation, data filtering, and agent tuning, leading to better generalization.
(2) We construct the GUI-Net-1M dataset, containing 1M data across five operating systems and over 280 applications. We have fully opened-source the code, the GUI-Net-1M dataset, and the trained models.
(3) We develop the TongUI agent by fine-tuning Qwen2.5-VL models on GUI-Net-1M, achieving improvements on multiple popular benchmarks.
2 Related Work
2.1 GUI Agent
The advancements of LLMs and VLMs accelerate the development of GUI agents. In the early state, only closed-source models are used (e.g., GPT-4, claude 3.5, and ChatGLM). Recently, more and more open-source models are released, such as Show-UI [18], UI-TARS [28], and CogVLM [38]. Grounding and planning are two important capabilities in GUI agents. The grounding capability means whether the GUI could identify correct buttons or regions for operations. Some methods improve the grounding capability by adding more prompts (accessibility trees, HTML, and set-of-mask) [49, 5, 8, 47, 23] and collecting grounding data for fine-tuning [43]. As for the planning capability of GUI agents, existing efforts improve it by prompt engineering [12, 1], supervised fine-tuning [18, 42, 10, 28], or reinforcement learning [13, 26]. In addition, some efforts focus on evaluating the GUI agents from multiple aspects, including action sequence [32, 9], grounding precision [4, 19], trajectory effectiveness [17, 16], and task completion [30, 35, 40]. The evaluating performance in turn guides the research of intelligent agents.
2.2 Agent Data Collection
Compared with data collection for LLMs or VLMs, collecting high-quality data for agents is more challenging, since agents usually require long trajectories that solve complex tasks in different domains (e.g., GUI, multimodal reasoning, embodied AI). Commonly used schemes include human annotation and model synthesis. Human annotation could produce high-quality and diverse trajectory data, but may result in high cost [20]. Some efforts have been made to use zero-shot agents to produce synthetic trajectory data [27, 21, 6], while this may cause out-of-distribution data, leading to inferior results. Similar to other domains, human annotation and model generation are also commonly used to collect trajectories for GUI agents, including the planning, action sequence, and action position [5, 32, 3, 28, 4, 22]. Note that the recent state-of-the-art method UI-TARS [28] has not release its data. In contrast, we will fully release our data.
In addition to the two schemes, considering that there are abundant resources for GUI operations on the Internet, some efforts focus on collecting trajectory data of GUI agents from the Web, to preserve the data quality and reduce costs simultaneously [25, 41]. They collect textual tutorials from the Internet, and convert them into trajectories by using LLMs or exploring in simulators. Unlike them, we directly collect multimodal tutorials from the Internet and propose a multimodal tutorial process pipeline to convert them into trajectories without any simulators. Meanwhile, compared with synthetic data in simulators, our data is closer to tasks in the real world with practical purposes, benefiting to the generalization capability of GUI agents.
Figure 2: Illustration for TongUI model architecture.
3 Method
3.1 Formulation
We formulate GUI tasks as a sequential decision process, and we adopt ReAct [44] as our agent framework. At one time step $i$, given the observation of previous $n$ steps $o_{i-n},\cdots,o_{i-1},o_{i}$ (e.g., a screenshot of GUI), the thoughts $r$ and actions $a$ of previous $n$ steps $(r_{i-n},a_{i-n},\cdots,r_{i-1},a_{i-1})$, and the query $q$, the agent generates a new thought $r_{i}$ and an executable action $a_{i}$ from the action space, such as clicking on a specific UI element, entering text, or scrolling through the interface. Then, executing the action $a_{i}$ leads to a new observation $o_{i+1}$ (e.g., an updated screenshot). This interaction loop continues, with the agent repeatedly observing the environment, selecting actions, and receiving updated observations until either a termination condition is met (e.g., the task is completed or fails) or the maximum number of steps is reached. We parameterize the GUI agent using a vision-language-action model $M_{\theta}$ as shown in Fig. 2,
$r_{i}^{\star},a_{i}^{\star}=\arg\max_{r_{i},a_{i}}M_{\theta}(r_{i},a_{i}|q,o_{i-n},r_{i-n},a_{i-n}\cdots,o_{i-1},r_{i-1},a_{i-1},o_{i}).$ (1)
In this paper, we propose the TongUI framework that aims to tune $M_{\theta}$ by learning from multimodal web tutorials, as shown in Fig. 3. Such tutorials are usually formatted as $\{v,e\}$, where $v$ denotes the visual information (such as images or videos), and $e$ denotes the textual information (such as an introduction to a video). Our goal is to collect diverse multimodal tutorials $\{v,e\}$, and convert them into training data for GUI agents $\{q,o_{1},r_{1},a_{1},\cdots,o_{T},r_{T},a_{T}\}$ with the query $q$ and the trajectory $(o_{1},r_{1},a_{1},\cdots,o_{T},r_{T},a_{T})$ in $T$ steps.
3.2 Tutorial Source
To collect multimodal web tutorials, we carefully select multiple data sources via brainstorming, ensuring that we cover a wide range of applications, operating systems, and task types. Concretely, we choose YouTube, Bilibili, TikTok, Wikihow, and Baidu Experience, which host a variety of user-generated content on GUI tasks. YouTube, Bilibili, and TikTok host a variety of video tutorials about GUI tasks. For video tutorials, the visual information $v$ is the video itself, and the textual information $e$ includes the title, the brief introduction to the video, and the caption or audio of the video. Wikihow and Baidu Experience are two widely used platforms that provide step-by-step articles with images in each step across multiple applications and sperating systems. Here, the visual information $v$ is the image sequence among all steps, and the textual information $e$ includes the title and textual content in the article. In this case, we could get diverse multimodal web tutorials.
Figure 3: Illustration of the TongUI framework.
3.3 Tutorial Crawling
We use a keyword-based search approach. The search keywords are constructed by combining the name of the applications or website with the task content, that is “app/web + task.” Here, “app/web” refers to the operating system or application being used (e.g., a mobile app or website), while “task” refers to a complete workflow or a set of actions required to accomplish a specific objective (e.g., changing font size in Word, or browsering in Chrome). We identify some task seeds across multiple applications via brainstorming, and expanded them using LLMs for more diverse and relevant tasks.
With these keywords, we crawl multimodal web tutorials from source websites. We employ platform-specific methods to retrieve tutorials, including subtitles and audio. For YouTube, we utilize the Google API and YouTube’s official API to search for videos, and download the relevant content. For Bilibili and TikTok, we use the unofficial API to search for videos and retrieve both video and audio streams for download. For Wikihow and Baidu Experience, articles usually have multiple tags to indicate the main category of the tutorial, such as Windows, Chrome, Word, and etc. We crawl text-image tutorials using tags.
3.4 Tutorial Process
Textual processing. Given a crawled multimodal tutorial $\{v,e\}$, this process extracts the task $q$ and rough descriptions $\{h_{1},\cdots,h_{T}\}$ in $T$ steps. In video tutorials from Youtube, Tiktok, and Bilibili, the task guidance is usually in the video’s audio information without any textual introductions. To solve this issue, we begin by applying the open-source speech recognition model, Whisper [29], to transcribe the audio streams into the text information $e$. For image-text articles on Wikihow and Baidu Experience, we parse websites based on their article structures for the textual information $e$.
We use LLMs to identify key task-related verbs and nouns, such as “clicking a button”, “filling out a form”, or “selecting a menu”. Then, we leverage both the title and extracted text to classify tutorials into three categories: mobile, desktop, and others. The data with category “others” is discarded. The categories “mobile” and “desktop” allow us to organize the data by device types, ensuring the tasks are appropriately tailored to their respective environments. Finally, we use LLMs to extract the task $q$ and summarize this process into rough descriptions $\{h_{1},\cdots,h_{T}\}$ of $T$ steps from the filtered textual information.
Visual processing. This process extracts the observation $\{o_{1},\cdots,o_{T}\}$ in the $T$ steps. For text-image articles, we could directly obtain the image sequences as the observation $\{o_{1},\cdots,o_{T}\}$ from the crawled data. In practice, there are some noisy data in Baidu Experience and Wikihow, which might be diagrams, comics, or natural images rather than GUI screenshots. To address this issue, we prompt GPT-4o-mini for image classifications, which classify each image as a screenshot or not.
For video tutorials, we need to extract key frames that correspond to critical actions, representing meaningful moments in the task-solving process. We observe that the audio transcript of the videos often contains valuable planning information, such as the description of task steps. Therefore, when audio is available, we first segment the video based on the audio transcript timestamps, treating each segment as a potential task step. We then extract key frames from each of these segments. In the absence of audio, we treat the entire video as a single segment and extract key frames throughout the video. In doing so, we adopt the MOG2 background subtraction algorithm [50, 51, 52] to detect significant changes in GUI interactions.
Figure 4: Data statistics on the GUI-Net-1M dataset.
After textual and visual processing, tutorials from different sources are structured into a task $q$ with text and image pairs of $T$ steps, denoted as ($\big{(}\{o_{1},h_{1}\},\cdots,\{o_{i},h_{i}\},\cdots,\{o_{T},h_{T}\}\big{)}$, where $o_{i}$ is the image and $h_{i}$ is the rough description in the $i$-th step, respectively.
3.5 Trajectory Generation
This process aims to generate the trajectory $(o_{1},r_{1},a_{1},\cdots,o_{T},r_{T},a_{T})$ in $T$ steps using the task $q$, and image $o_{i}$ and rough description $h_{i}$ in $T$ steps. Concretely, we use a pretrained GUI agent (such as UI-TARS [28]) to generate the thought $r_{i}$ and the action $a_{i}$ using the same action space defined by the GUI agent. For the thought $r_{i}$ and action $a_{i}$, we feed the observation $o_{i}$ with $h_{i}$ as the query to the zero-shot agent. Here, we use $h_{i}$ as the query for agents instead of the task $q$, because $q$ usually contains an abstract goal instead of specific instruction about what to do on the screen. We empirically find that using $h_{i}$ leads to better performance. In this case, each step of text and image pairs can be written as $\{o_{i},r_{i},a_{i}\}$. In practice, the model might fail to generate well-formatted actions. We discard this step and make the next action the beginning of a new trajectory. For example, for a 4-step trajectory which has a failed generation in the second step, we split it into two training trajectories: ($(\{o_{1},r_{1},a_{1}\})$ and ($(\{o_{3},r_{3},a_{3}\},\{o_{4},r_{4},a_{4}\})$. Finally, we combine the trajectory with the task $q$ as $(q,\{o_{i},r_{i},a_{i}\}_{i=1}^{T})$. We collect these trajectories into a GUI-Net-1M dataset that contains 1M data. These data is on five operating systems with more than 280 applications. After that, we use GUI-Net-1M for supervised fine-tuning (SFT).
3.6 Data Filtering
We apply a multi-stage data filtering pipeline for data quality, as shown in Fig. 3. (1) Duplicate Tutorial Removal. To eliminate redundancy of crawling, we directly remove exact duplicate tutorials based on their unique identifiers (e.g., video IDs and URLs). This step ensures that each retained tutorial is unique, avoiding repetition and simplifying downstream processing. (2) Content Filtering. After tutorial processing, we employ an LLM to semantically filter out tutorials that are irrelevant to the GUI tasks. Concretely, based on the tutorial titles and audio transcriptions, the LLM evaluates whether the content is centered around GUI interactions. (3) Trajectory Filtering. In the trajectory generation stage, since we use UI-TARS [28], if a screenshot corresponds to a step unrelated to GUI interaction, the agent typically predicts actions like wait or call_user. We use this as a signal to discard such observations. Then, we feed the screenshot and trajectory of a task into a Qwen2.5-VL-7B model, and prompt it to filter out low-quality data.
3.7 Agent Tuning
Given a data point $(q,\{o_{i},r_{i},a_{i}\}_{i=1}^{T})$ of $T$-step, we train a VLM $M_{\theta}$ using the cross-entropy loss,
$\min\mathbb{E}_{(q,\{o_{i},r_{i},a_{i}\}_{i=1}^{T})\sim\mathbb{D}}\big{[}-\sum_{i=1}^{T}M_{\theta}(r_{i},a_{i}|q,o_{i-n},r_{i-n},a_{i-n},\cdots,o_{i-1},r_{i-1},a_{i-1},o_{i})\big{]},$
where $\mathbb{D}$ is the collected GUI-Net-1M dataset, and we sum the loss values of the $T$ steps in the trajectory. After training, we obtain the TongUI agent.
4 Dataset Statistics
We provide four key statistics to show the diversity of the collected GUI-Net-1M dataset. We show the operating system distribution in  Fig. 4(a). Our dataset covers a diverse range of operating systems, including Windows, Android, iOS, MacOS, and Linux, ensuring a broad representation of GUI interactions across both desktop and mobile environments. The application distribution is shown in  Fig. 4(b). Our dataset covers a wide variety of application categories, ensuring a broad and diverse range of user tasks. We show the step distribution in  Fig. 4(c). Our dataset includes GUI interactions with varying step lengths, ranging from single-step actions to 9-step tasks. The distribution shows that shorter tasks (1-3 steps) are more frequent, while longer tasks gradually decrease in numbers. The higher proportion of trajectories of 1-3 steps is partly due to the trajectory splitting strategy described in Section 3.5. The action distribution is shown in  Fig. 4(d). Click and Tap are the most frequent, reflecting their central role in both desktop and mobile interactions. Other common actions include HotKey, Swipe, and Drag, covering various forms of user interactions.
Figure 5: Data flow of data filtering.
We also provide statistics on the data filtering flow in  Fig. 5. Throughout the data filtering pipeline, we observe a gradual reduction in data volume, reflecting the progressive refinement of the dataset. After Duplicate Tutorial Removal, 72.3% of the data remains. Following Content Filtering, this is further reduced to 58.5%. Finally, after applying Trajectory Filtering, 33.0% of the original data is retained. This step-wise filtering process helps to include high-quality and GUI-relevant interactions in the final dataset.
Model Desktop Icon Desktop Text Mobile Icon Mobile Text Web Icon Web Text Average
SeeClick-9.6B [4] 30.0 72.2 52.0 78.0 32.5 55.7 53.4
UGround-7B [7] 63.6 82.5 60.3 82.8 80.4 73.3 70.4
OmniParser-GPT-4V [23] 63.6 91.3 57.0 93.9 51.0 81.3 73.0
ShowUI-2B [18] 61.1 76.3 75.5 92.3 63.6 81.7 75.1
Qwen2.5-VL-3B $\dagger$ 30.7 66.0 40.2 70.0 51.5 69.1 56.5
Qwen2.5-VL-7B $\dagger$ 60.0 86.6 71.6 91.2 67.5 85.2 78.6
TongUI-3B 77.1 92.3 77.7 92.6 74.8 87.8 83.6
TongUI-7B 80.0 93.8 79.5 91.9 81.6 89.1 86.0
Table 1: Results on ScreenSpot; $\dagger$ means the results are re-produced. The best method is marked in bold, and the second-best method is underlined.
Model Mobile Desktop Web Avg
Text Icon/Widget Text Icon/Widget Text Icon/Widget
SeeClick [4] 78.4 50.7 70.1 29.3 55.2 32.5 55.1
OS-Atlas-4B [39] 87.2 59.7 72.7 46.4 85.9 63.1 71.9
OS-Atlas-7B [39] 95.2 75.8 90.7 63.6 90.6 77.3 84.1
Qwen2.5-VL-3B $\dagger$ 77.2 54.5 74.7 35.7 68.8 60.1 64.2
Qwen2.5-VL-7B $\dagger$ 95.5 81.5 90.2 65.0 90.2 70.4 84.0
TongUI-3B 94.4 79.6 92.8 75.0 87.6 77.8 85.5
TongUI-7B 93.1 81.5 96.4 82.9 90.2 84.7 88.7
Table 2: Results on ScreenSpot-V2.
Agent Model Development Creative CAD Scientific Office OS Avg
Text Icon Avg Text Icon Avg Text Icon Avg Text Icon Avg Text Icon Avg Text Icon Avg Text Icon Avg
GPT-4o 1.3 0.0 0.7 1.0 0.0 0.6 2.0 0.0 1.5 2.1 0.0 1.2 1.1 0.0 0.9 0.0 0.0 0.0 1.3 0.0 0.8
SeeClick [4] 0.6 0.0 0.3 1.0 0.0 0.6 2.5 0.0 1.9 3.5 0.0 2.0 1.1 0.0 0.9 2.8 0.0 1.5 1.8 0.0 1.1
OS-Atlas-4B [39] 7.1 0.0 3.7 3.0 1.4 2.3 2.0 0.0 1.5 9.0 5.5 7.5 5.1 3.8 4.8 5.6 0.0 3.1 5.0 1.7 3.7
ShowUI-2B [18] 16.9 1.4 9.4 9.1 0.0 5.3 2.5 0.0 1.9 13.2 7.3 10.6 15.3 7.5 13.5 10.3 2.2 6.6 10.8 2.6 7.7
CogAgent-18B [10] 14.9 0.7 8.0 9.6 0.0 5.6 7.1 3.1 6.1 22.2 1.8 13.4 13.0 0.0 10.0 5.6 0.0 3.1 12.0 0.8 7.7
Aria-UI [43] 16.2 0.0 8.4 23.7 2.1 14.7 7.6 1.6 6.1 27.1 6.4 18.1 20.3 1.9 16.1 4.7 0.0 2.6 17.1 2.0 11.3
UGround-7B [7] 26.6 2.1 14.7 27.3 2.8 17.0 14.2 1.6 11.1 31.9 2.7 19.3 31.6 11.3 27.0 17.8 0.0 9.7 25.0 2.8 16.5
OS-Atlas-7B [39] 33.1 1.4 17.7 28.8 2.8 17.9 12.2 4.7 10.3 37.5 7.3 24.4 33.9 5.7 27.4 27.1 4.5 16.8 28.1 4.0 18.9
Qwen2.5-VL-3B $\dagger$ 14.9 2.1 8.7 7.1 2.1 5.0 1.5 1.6 1.5 13.9 8.2 11.4 10.7 3.8 9.1 9.4 2.3 6.1 9.1 3.3 6.9
Qwen2.5-VL-7B $\dagger$ 22.1 0.7 11.7 16.2 4.2 11.1 5.1 3.1 4.6 25.0 7.3 17.3 22.0 11.3 19.6 17.8 4.5 11.7 17.4 4.5 12.5
TongUI-3B 32.5 0.7 17.1 24.8 2.8 15.5 11.7 1.6 9.2 43.1 12.7 29.9 32.8 7.6 27.0 15.0 1.1 8.7 26.4 4.1 18.0
TongUI-7B 40.9 3.5 22.7 31.3 7.0 21.1 17.3 9.4 15.3 50.7 12.7 34.3 45.8 13.2 38.3 28.0 6.7 18.4 35.1 8.0 24.7
Table 3: Results on ScreenSpot-Pro.
Method General Single Web Shopping Install Google Apps Average
PaLM2-CoT [48] 39.6
OmniParser [23] 48.3 57.8 51.6 77.4 52.9 57.7
SeeClick-9.6B [4] 54.0 73.7 57.6 66.4 54.9 59.3
ShowUI-2B [18] 63.9 77.5 66.6 72.5 69.7 70.0
[…]
Qwen2.5-VL-3B-ShowUI 66.0 74.4 65.0 74.5 70.3 70.1
TongUI-3B 65.6 77.0 65.8 75.1 74.5 71.6
TongUI-7B 67.6 79.9 69.1 76.3 73.5 73.3
Table 4: Results on the AITW. We report results on five splits of AITW and the average scores.
5 Experiments
5.1 Setting
Additional Training Data. In addition to GUI-Net-1M, we also refine some publicly available GUI datasets. For GUI grounding, we adopt ShowUI-website [18], ShowUI-desktop [18], and AMEX [2]. For GUI navigation, following the guidance of ShowUI, we use the training set from AITW [31], Mind2Web [5], MiniWob [11], and GUIAct [3]. We prompt GPT-4o-mini to generate thoughts for these datasets with textual guidance and visual spots on images, obtaining 256K data. We combine these data with GUI-Net-1M for training, totally 1M+256K $\approx$ 1.3M data. To make sure the thought is consistent with the action, we explicitly add vision prompts to images, such as a spot that indicates the location of the clicked button. In our text prompt, we instruct the model to generate a thought based on textual guidance of the task, description of the action, and vision prompts on the image.
Evaluation Data. We evaluate the TongUI agent on offline benchmarks: ScreenSpot [4], ScreenSpot-V2 [39], ScreenSpot-Pro [15], AITW [31], AndroidControl [16], Mind2Web [5], and Baidu Experience (we manually annotate 102 data points from the collected Baidu Experience data for testing). We also evaluate the TongUI agent on an online benchmark: MiniWob [34]. We use the same 35 tasks selected by SeeClick[4] to evaluate our model.
Training. We use Qwen2.5-VL-3B/7B as our base vision-language-action model. We train our model with a context window of 8192. We set the maximum number of previous observations as 2, which results in at most three images in a trajectory. We limit each image to 1350 vision tokens. Our training involves two stages. Following existing methods [18], we utilize the SFT data (training set) of the evaluation dataset to fine-tune the model. We apply LoRA on our model’s vision encoder and LLM parts with rank 16 and alpha 32. This setting leads to 0.5% of total parameters. We set the batch size per device to 4 and the gradient accumulation steps to 8. We set the learning rate as $41e-$ 4 and AdamW optimizer for training. The training process on 8$\times$×A100-80G takes roughly four days for a 3B model and seven days for a 7B model.
Compared Methods. We mainly compare the TongUI agent with open-source methods, including ShowUI [18], SeeClick [4], OS-Atlas [39], CogAgent [10], etc.
Additional experimental results, such as data scaling effect, are presented in the Supplementary Material.
5.2 Main Results
5.2.1 Grounding
In Tab. 1,  Tab. 2, and  Tab. 3, we show the zero-shot grounding performance of TongUI on ScreenSpot, ScreenSpot-V2, and ScreenSpot-pro, respectively. Similar to previous works [18], grounding on icon is much harder than grounding on text. The collected data leads to significant improvements on the baseline Qwen2.5-VL model. Compared with ShowUI, our method has about 5 % - 20 % improvements. These results demonstrate that GUI-Net-1M can indeed improve the grounding capability of GUI agents.
5.2.2 Offline Navigation
AITW. AITW is a dataset about navigation on Android. The comparisons are shown in Tab. 4. We observe that the fine-tuned TongUI-3B model achieves better performance compared to ShowUI-2B by 1.6% in overall success rate. This shows that the collected data also improves the reasoning and planning capabilities of GUI agents on the navigation tasks. The larger model, TongUI-7B, gains 1.7% improvement compared to TongUI-3B, which is consistent with common sense.
Method Model High Low
AITW PaLM 2L 19.5 56.7
SeeClick GPT-4-Turbo 33.9 54.3
M3A GPT-4-Turbo 42.1 55.0
ER PALM-2S 19.5 45.5
ER PALM 2L 33.0 45.9
ER GPT-4 32.1 51.7
ER Gemini 1.5 Pro 24.4 50.2
TongUI TongUI-3B 32.0 52.0
TongUI TongUI-7B 38.2 69.3
Table 5: Step accuracy on AndroidControl.
AndroidControl. We also evaluate our performance in the AndroidControl benchmark in Tab. 5. As we do not include the AndroidControl training set, we compare our model with the zero-shot methods mentioned in previous work  [16]. TongUI-7B model achieves the best performance of low-level instructions(12.6% improvement).
Mind2Web. In Tab. 6, the results of web navigation show that TongUI achieves the best performance on all metrics. Notably, cross-task, cross-website, and cross-domain tasks are more difficult than conventional GUI tasks. This highlights that the collected data improves the generalization capability of GUI agents. We argue that the reason is that the collected data involves diverse applications and operating systems, improving the generalization capability. Compared to ShowUI-2B, TongUI-3B has obvious improvements, especially on Elem. Acc and Step SR. Compared with ShowUI-Qwen2.5-VL-3B that fine-tunes the Qwen2.5-VL-3B model using the ShowUI dataset, TongUI-3B also has better performance (3.2% improvements), showing the quality of our data.
Method Cross-Task Cross-Website Cross-Domain
Elem. Acc OP. F1 Step SR Elem. Acc OP. F1 Step SR Elem. Acc OP. F1 Step SR
CogAgent [10] 22.4 53.0 17.6 18.4 42.4 13.4 20.6 42.0 15.5
MindAct [5] 55.1 75.7 52.0 42.0 65.2 38.9 42.1 66.5 39.6
OmniParser [23] 42.4 87.6 39.4 41.0 84.8 36.5 45.5 85.7 42.0
ShowUI-2B [18] 39.9 88.6 37.2 41.6 83.5 35.1 39.4 86.8 35.2
SeeClick-9.6B [4] 28.3 87.0 25.5 21.4 80.6 16.4 23.2 84.8 20.8
Qwen2.5-VL-3B $\dagger$ 2.5 14.5 0.4 2.7 12.6 1.0 3.3 24.2 1.7
Qwen2.5-VL-7B $\dagger$ 6.2 72.8 5.0 6.3 68.2 4.5 8.4 73.6 7.2
Qwen2.5-VL-3B-ShowUI 43.2 88.7 39.7 41.3 86.7 35.5 45.1 86.1 40.7
TongUI-3B 53.4 89.0 48.8 54.2 86.4 48.1 53.8 88.2 49.5
TongUI-7B 58.1 88.7 53.4 55.6 87.2 49.0 57.6 88.7 52.9
Table 6: Results on Mind2Web. We report results on three types of tasks: cross-task, cross-website, and cross-domain. “Elem. Acc” means whether the element is selected correctly, “OP. F1” denotes the F1 score for the predicted action, and “Step SR” counts successful steps.
Baidu Experience. The collected Baidu experience evaluation set includes tasks in Chinese, which evaluate the performance of GUI agents across different languages. Results are shown in Tab. 7. Compared with the baseline model, TongUI has improvements on both action accuracy and success rate, showing better generalization capability brought by the GUI-Net-1M dataset.
Model Action Acc. Success
ShowUI-2B [18] 65.7 14.7
Qwen2.5-VL-3B 36.3 10.8
Qwen2.5-VL-7B 70.6 16.6
TongUI-3B 86.3 18.6
TongUI-7B 87.3 25.5
Table 7: Results on Baidu Experience.
5.2.3 Online Navigation
We evaluate the online navigation performance on the MiniWob benchmark, and the results are shown in Tab. 8. Compared with offline navigation, online navigation tasks are more challenging, since they require real-time processing of dynamic environments, handling unexpected obstacles, and adapting to changes in the navigation path. In this case, TongGUI that learns from multimodal tutorials achieves improvements again. Considering that the multimodal tutorials are offline data, this improvements confirm the generalization of TongUI.
Model FT Score
CC-Net (SL) [11] $\checkmark$ 23.4
Pix2Act [33] $\checkmark$ 55.2
SeeClick-9.6B [4] $\checkmark$ 67.0
Qwen2-VL-2B [36] $\checkmark$ 66.8
ShowUI-2B [18] $\checkmark$ 71.5
Qwen2.5-VL-3B $\times$ 0.3
Qwen2.5-VL-3B-ShowUI $\checkmark$ 67.7
TongUI-3B $\checkmark$ 72.7
TongUI-7B $\checkmark$ 73.9
Table 8: Results on MiniWob.
5.3 User Study for Data Quality
To validate the effectiveness of our data filtering strategy and evaluate the data quality of the collected GUI-Net-1M dataset, we conduct two user studies. In both studies, five participants with substantial experience in GUI agent research (but not involved in our project) are asked to rate the quality of each data point on a scale from 0 (very poor) to 5 (excellent). As shown in Fig. 6(a), we evaluate the impact of our Trajectory Filtering step. Specifically, we randomly sample 100 data points before filtering and 100 after filtering, and mix them. As shown in Fig. 6(a), the average score increases from 3.22 to 4.12 after filtering, demonstrating that Trajectory Filtering significantly improves data quality by effectively removing low-quality or GUI-irrelevant steps. We also randomly sample 100 data points from the ShowUI dataset for comparison. The results are shown in Fig. 6(b), the ShowUI dataset received an average score of 4.26. This suggests that data quality in GUI-Net-1M is comparable to that in the ShowUI dataset, while our data leads to further improvements.
Figure 6: Average scores from humans
6 Conclusion
In this paper, we present the TongUI framework that allows GUI agents to learn from multimodal web tutorials. By defining suitable and rich sources, we can crawl multimodal tutorials about diverse applications in different operating systems. The proposed tutorial processing method can extract tasks and trajectories from tutorials, and we obtain the GUI-Net-1M dataset. Using this dataset, the tuned Qwen2.5-VL model achieves improvements on multiple commonly used benchmarks, demonstrating the effectiveness of the TongUI framework and the collected GUI-Net-1M dataset. In this method, we have to collect all tutotrials and train the model once, which may overlook the potential for continual learning. In the future, we will explore the continual learning capability of GUI agents, enabling the agents to better adapt to new environments.